OpenStudy (almostdone):

The perimeter of a television screen is 150 inches. The ratio of the length to the width is 16 to 9 find the length and width.

OpenStudy (3mar):

May I help?

OpenStudy (rcp031):

Only if television screen is Rectangular Then, Perimeter P=2L+2W HERE =LENGTH W=WIDTH the ratio of the length to the width is 16 to 9 ie, L:W=16/9 & P=150 Therefore L/W=16/9, or W/L=9/16 ie, W=16L/9 P=2L+2W =2L+2(16L/9)=150 DO THE MATH GET ANSWER!

OpenStudy (rcp031):

Sorry, typos it should be W=9L/16 P=2L+2W =150 2L+2W=150 2L+2(9L/16)=150

OpenStudy (almostdone):

Thanks